# Hierarchical Mixture of Experts: Generalizable Learning for High-Level Synthesis

### v1

@@ Abstract
High-level synthesis (HLS) is a widely used tool in designing Field Programmable Gate Array (FPGA). HLS enables FPGA design with software programming languages by compiling the source code into an FPGA circuit. The source code includes a program (called “kernel”) and several pragmas that instruct hardware synthesis, such as parallelization, pipeline, etc. While it is relatively easy for software developers to design the program, it heavily relies on hardware knowledge to design the pragmas, posing a big challenge for software developers. Recently, different machine learning algorithms, such as GNNs, have been proposed to automate the pragma design via performance prediction. However, when applying the trained model on new kernels, the significant domain shift often leads to unsatisfactory performance. We propose a more domain-generalizable model structure: a two-level hierarchical Mixture of Experts (MoE), that can be flexibly adapted to any GNN model. Different expert networks can learn to deal with different regions in the representation space, and they can utilize similar patterns between the old kernels and new kernels. In the low-level MoE, we apply MoE on three natural granularities of a program: node, basic block, and graph. The high-level MoE learns to aggregate the three granularities for the final decision. To stably train the hierarchical MoE, we further propose a two-stage training method. Extensive experiments verify the effectiveness of the hierarchical MoE.
@@ Introduction
With the heated demand for domain-specific accelerators, field-programmable gate arrays (FPGA) are widely used. It is, however, labor-intensive to write register-transfer-level hardware description languages, such as VHLD and Verilog. High-level synthesis (HLS) provides a much easier solution by compiling a source code written in a software programming language, such as C, MatLab, etc., into an FPGA circuit (Cong et al. 2011, 2022; Schafer and Wang 2019).
Figure 1: Motivation of utilizing MoE. The two kernels share a similar nested loop, while they are different in codes outside this screenshot. Similar blocks/graphs can utilize similar experts via similar gating, while different blocks/graphs can utilize different experts.
@@
While machine learning models have better learning ability than the heuristic methods, they do not generalize well to unseen kernels. In real-world applications, we often encounter new circuit design requirements. The model trained on existing kernels usually fails to perform well on new kernels. This can be viewed as a domain generalization (Bai et al. 2022; Kwon and Carloni 2020) problem, where each kernel is a domain. Note it is very time-consuming to run HLS to acquire the labels on new kernels, with each run taking minutes to hours (Sohrabizadeh et al. 2023), thus data-efficient fine-tuning is required for new kernels. Domain generalization methods for GNN usually employ adversarial training to align the representation space between the source domain and target domain (Dai et al. 2019; Zhang et al. 2019; Wu et al. 2020; Shen et al. 2023), or do data augmentation to learn invariant representation under risk extrapolation (Wu et al. 2022b; Liu et al. 2023a). However, these approaches are not applicable to our application, as the differences between kernels are informative for the prediction, so forcing them to the same distribution does not work.
@@
Nonetheless, a unique opportunity of the HLS prediction task is that kernels usually share some similar substructures, as shown in Fig. 1. It could be beneficial if we can leverage such similarities. We thus propose a two-level hierarchical mixture of experts (MoE) to address this issue. MoE (Jacobs et al. 1991; Shazeer et al. 2017) uses several expert networks instead of a single fixed network, and it computes the weighted sum of their outputs. Different experts can specialize in different regions in the hidden representation space, so that similar substructures in different domains can be processed by the same expert. A gating network computes the expert assignment, which learns to discern which expert should deal with which substructure. GraphMETRO (Wu et al. 2024) applies MoE to improve GNN’s generalizability by predefining a set of domain shifts and training each expert to deal with one type of domain shift by data augmentation. We take an orthogonal direction, eliminate the need for hardware knowledge to define the domain shifts, and focus on the model structure rather than data augmentation.
@@
While designing the model architecture is nontrivial, another unique opportunity for HLS prediction is that the input graph naturally has three granularities: nodes, blocks, and the whole graph. MoE can be constructed on either of these three granularities. Then, which granularity to choose? It is a hard task as each kernel might benefit from different granularities differently. Simply applying MoE on all three granularities in one model does not work. Therefore, we propose a higher-level MoE to aggregate the three low-level MoEs according to the need. Nonetheless, optimization with the hierarchical MoE is challenging. Since the three low-level MoEs are different, training easily leads to expert polarization. We propose a two-stage training and constant initialization to stabilize the training. To the best of our knowledge, while some papers (Shazeer et al. 2017) have studied hierarchical MoE to improve memory efficiency, this is the first study that demonstrates the performance gain of hierarchical MoE, and it might also potentially benefit other applications. We conduct extensive experiments on the largest HLS benchmark dataset, HLSyn (Bai et al. 2023), and the experiment results reveal the effectiveness of hierarchical MoE. In summary, we make the following contributions:
@@
  • We propose the hierarchical MoE to address these challenges, and a two-stage training to stabilize its training.
@@
There are two research directions for HLS pragma design automation. The first direction is heuristic-driven. AutoDSE (Sohrabizadeh et al. 2021) uses bottleneck-guided searching based on the feedback from the HLS report, and it is well-generalizable. Non-linear programming performs well on affine kernels based on a lower-bound objective function (Pouget, Pouchet, and Cong 2024a, b). The second direction is data-driven, which trains a surrogate machine learning model to predict the FPGA’s performance. GNNs (Sohrabizadeh et al. 2022, 2023; Ustun et al. 2020; Wu et al. 2022a; Wu, Xie, and Hao 2023, 2021; Qin et al. 2024) are widely used. While machine learning models could perform better than heuristic methods with sufficient training data, they suffer from poor generalizability. To solve this issue, previous works use meta-learning method, MAML (Finn, Abbeel, and Levine 2017), to find a generalizable parameter initialization (Bai et al. 2022), or train a separate domain adaptor for each domain (Kwon and Carloni 2020). They perform well in some situations, but in our case, we have too many kernels to stably train MAML, and we only have hundreds of data for each kernel, so it is difficult to train an individual adaptor. Instead, we explore an orthogonal direction of improving the generalizability of the model structure.
@@
Domain generalization aims to reduce the performance gap between the source domains and target domains. Most related works (Dai et al. 2019; Zhang et al. 2019; Wu et al. 2020; Shen et al. 2023) employ adversarial training to align the representation space between the source and target domains. Some papers (Wu et al. 2022b; Liu et al. 2023a) design specific data augmentation methods and learn invariant representations under risk extrapolation. However, we cannot directly align the representation space of different kernels since their difference is large and informative. GraphMETRO (Wu et al. 2024) predefines several types of domain shifts and uses MoE to deal with them, where each expert deals with one type of domain shift. While it is a good general method, we take an orthogonal direction and eliminate the need for extensive hardware knowledge to define the domain shifts, and our model based on the three granularities is specifically designed for the HLS prediction task.
@@
MoE trains several expert networks and calculates the weighted sum of their outputs (Jacobs et al. 1991; Shazeer et al. 2017). It is useful in domain transfer learning in computer vision (Li et al. 2023; Zhong et al. 2023). It has also been utilized in GNN to improve the performance (Wang et al. 2023; Han et al. 2024), diversify node representations in fairness-augmented graphs (Liu et al. 2023b), and deal with class imbalance (Hu et al. 2021). Previous work (Shazeer et al. 2017) has employed hierarchical expert routing to improve memory efficiency, enabling a larger number of experts. However, previous papers have not shown the performance gain of hierarchical MoEs than regular MoEs when using the same number of total experts.
@@
HLS prediction is formalized as a graph regression task. Following previous works (Sohrabizadeh et al. 2022, 2023), we utilize ProGraML (Cummins et al. 2021) graph to represent a source code, which is a control data flow graph. Nodes represent instructions, variables, and constant values, and edges represent the control flow and data flow. A GNN model is trained to predict the FPGA’s latency and the utilization of four resources: LUT, FF, DSP, and BRAM. In the domain generalization setting, we train the model on $N$ source kernels $D^{(train)}=\{D_{1},D_{2},...,D_{N}\}$, where $D_{i}=\{(X_{i},T_{i1},Y_{i1}),(X_{i},T_{i2},Y_{i2}),...,(X_{i},T_{ii_{n}},Y_{ii_{n}})\}$ is the i-th kernel containing $i_{n}$ samples, and each sample consists of a program $X$, a pragma design $T$, and a label $Y$. For a new kernel $D_{test}$, we mimic the data-scarce situation where we only use $k$ samples from the dataset $D^{(k)}\subset D_{test},|D^{(k)}|=k$, to fine-tune the model. After fine-tuning, there are two ways of evaluation: offline evaluation and online evaluation. In the offline evaluation, we evaluate the MSE on the left-out test samples: $D_{test}\setminus D^{(k)}$. In the online evaluation, we run a DFS searching algorithm used by previous works (Sohrabizadeh et al. 2022, 2023) to search for good pragma designs. Based on the fine-tuned model’s prediction, We select the top-10 pragma designs, run HLS for them, and evaluate the FPGA’s speedup compared to AutoDSE.
@@
The proposed hierarchical MoE can be adapted to any GNN model. We use one of the SOTA GNN models for HLS prediction, HARP (Sohrabizadeh et al. 2023), as the base model. Since the original ProGraML graph is sparse and not good at modeling long-term dependency, HARP creates a “pseudo node” for each basic block in the program and connects it with every node within that basic block. Defined by the compiler community, a basic block is a sequence of instructions that has a single entry point and a single exit point, whose terminator instruction can be branch, return, etc. We illustrate HARP’s model structure in the appendix. We use $V$ to denote the set of all nodes and $V_{B}$ to denote the set of pseudo nodes. HARP consists of five components: (1) GNN encoder: it consists of 6 GNN layers and learns node representation; (2) Pragma MLP: it uses an MLP for each pragma type to transform the representation of pseudo nodes that are modified by that pragma type; (3) Another GNN layer after pragma MLP: it spreads the information of pragmas from pseudo nodes to normal nodes and to pseudo nodes that do not have pragmas; (4) Graph pooling: it does graph pooling on pseudo nodes $V_{B}$ to form a graph representation and on normal nodes $V\setminus V_{B}$ to form another graph representation, then it concatenates them; (5) Output MLP: it makes the prediction based on the graph representation.
@@ Methods
A unique opportunity for our task is that the input graph has three granularities: normal nodes that represent data and instructions, pseudo nodes that represent basic blocks, and the whole graph that represents a source code file. If a data point from the target kernel shares a similar statement (node), basic block, or even the whole code (graph) with a data point from the training kernels, MoE could be helpful. Thus, we consider employing MoE on the three granularities: MoE on all nodes including normal nodes and pseudo node (node MoE), MoE only on pseudo nodes (block MoE), and MoE on the graph (graph MoE). Based on our pre-exploration, the best practice is to apply MoE in the components after the pragma MLP, since the experts need to share the same encoder and pragma MLPs. Our design of the three low-level MoE models is illustrated in Figure 2.
@@
$\boldsymbol{h_{i}^{\prime}}=\sum_{j=1}^{n}softmax(\boldsymbol{W_{G}}\boldsymbol{h_{i}})_{j}\cdot GNN_{j}(\boldsymbol{h_{i}},Graph),$ (1)
where $\boldsymbol{W_{G}}\in\mathbb{R}^{n\times d}$ is the parameter of the gating network, and $GNN_{j}(\cdot)$ is the j-th expert. The $softmax$ function ensures that all experts’ weights sum up to 1. By employing the MoE layer, different nodes can utilize different parameters for message passing and message aggregation.
Block MoE. We have tried applying MoE on the pragma MLP, but the performance has decreased. It might be because each pragma type has a separate pragma MLP, so the training data for each pragma MLP is already not abundant. If we further apply MoE, it will dilute the training data even more and hinder the performance. Instead, we add an additional MoE layer on the pseudo nodes’ representations after the pragma MLP, so that we can increase the expressiveness of the pseudo nodes’ representations before graph pooling. The representation space of the pseudo nodes after the block MoE could be different from normal nodes, so it is no longer suitable to have another GNN layer after the pragma MLP to do message passing between pseudo nodes and normal nodes. We have also tried applying an additional layer for normal nodes after the pragma MLP, but the performance is unsatisfactory. It might be because the normal nodes lack a global view, so it is not useful to further transform its representation by another layer. After having the block MoE, doing graph pooling only on the pseudo nodes performs better than doing graph pooling on both pseudo nodes and normal nodes. Therefore, we do not utilize normal nodes after the GNN encoder in the block MoE model.
@@
$\boldsymbol{h_{i}^{\prime}}=\sum_{j=1}^{n}softmax(\boldsymbol{W_{G}}\boldsymbol{h_{i}})_{j}\cdot\boldsymbol{W_{ej}}\boldsymbol{h_{i}}\ \ (v_{i}\in V_{B}),$ (2)
where $\boldsymbol{W_{G}}\in\mathbb{R}^{n\times d}$ is the parameter of the gating network, and $\boldsymbol{W_{ej}}\in\mathbb{R}^{d\times d}$ is the j-th expert network.
@@
$\hat{Y}^{(t)}=\sum_{j=1}^{n}softmax(\boldsymbol{W_{G}}\boldsymbol{h_{G}})_{j}\cdot MLP^{(t)}j(\boldsymbol{h_{G}}),$ (3)
where $MLP^{(t)}_{j}(\cdot)$ is the j-th expert MLP for predicting the target $t$ (the target could be the FPGA’s latency or a certain resource’s utilization). It contains four linear layers and the ELU activation function between them. The graph MoE enables different design points to utilize different parameters to make the final prediction.
Regularization term. As discovered in many literature (Shazeer et al. 2017; Wang et al. 2023; Li et al. 2023), expert polarization is a common problem of MoE. Due to the random initialization of expert networks, different experts initially have different performances, and the gating network learns to assign higher weights to better experts. It results in more training of better experts, leading to their better performance, which in turn leads to their higher weights. As training continues, the MoE model might collapse and only use the best expert. To encourage balanced expert assignments, we apply a regularization term (Shazeer et al. 2017):
@@
where $CV(\cdot)$ is the coefficient of variation. We calculate the importance score $\boldsymbol{I(W_{G})}\in\mathbb{R}^{n}$ for gating network $W_{G}$ as the total weights assigned to each expert. For node MoE, $M$ is the total number of nodes in all graphs; for block MoE, $M$ is the total number of pseudo nodes in all graphs; for graph MoE, $M$ is the number of graphs.
@@
We propose two designs of the high-level gating network, as illustrated in Figure 2. The first design is to do graph pooling on the input node features to form a graph representation as the input to the high-level gating network. We use the self-attention graph pooling. Denoting the input feature of node $v_{i}$ as $\boldsymbol{x_{i}}$, the graph pooling is formalized as:
@@
where $X_{G}$ is the aggregated input feature, and $V$ is the set of all nodes. The second design is to concatenate the graph representation in the three low-level MoE models. The second design performs better in our experiments, since the hidden representations are more expressive than the input features. Nonetheless, when we use a sparse MoE where only one or two experts are selected, the first design will be more memory efficient. It can determine the expert assignment before the computation of three expert models, thus reducing unnecessary computation. However, the best-performing method is to utilize all the experts, and in this case, the two designs are similar in efficiency. We utilize all the experts in our experiments. We also apply the regularization term for the high-level gating network.
Two-Stage Training and Constant Initialization
Optimizing the hierarchical MoE is challenging. Different from previous MoE studies where expert networks have the same or similar structures, our three low-level MoEs are very different and thus have different convergence speeds. The graph MoE model converges the fastest, since its MoE operates on the graph representation and has the least computation. Thus, the high-level gating network suffers severely from expert polarization, leaning towards the graph MoE model. If we simply increase the weight of the regularization term, the high-level gating network can effectively learn to assign about $\frac{1}{3}$ weight to each expert, but the graph MoE model still converges much faster than the other two models. In the first few epochs, the output of node MoE and block MoE models will be noisy and thus they are discouraged from making predictions. As a result, the graph MoE model will learn to output three times the label, while node and block MoE models will learn to output zero.
To address this unique challenge, we design a two-stage training strategy to encourage every expert model to perform well individually. In the first stage containing $T$ epochs (warmup), we train the three expert models individually. In the second stage, we take turns training the whole model end-to-end and the three expert models individually. If we denote the label as $Y$, the prediction made by the $i$-th expert model as $\hat{Y}_{i}$, and the MSE loss function as $L(Y,\hat{Y_{i}})$, then we define the loss function at epoch $t$ as:
@@
Here, $L_{R}=\frac{1}{3}(L_{R1}+L_{R2}+L_{R3})$, where $L_{Ri}$ is the regularization term of the i-th low-level MoE. $L_{Rh}$ is the regularization term of the high-level MoE. $g_{i}$ is the weight assigned by the high-level gating network. When we fine-tune our model on target kernels, we directly use the end-to-end joint training, since all the experts can already perform well during fine-tuning and there is no risk of expert polarization.
In common situations where the expert networks have similar or the same structures, the gating network’s parameters should be randomly initialized, because we need the randomness to diversify the experts. Instead, we initialize the high-level gating network to assign the same weights to the expert models, which can further prevent expert polarization. For low-level MoEs, we use random initialization.

### v2

@@ Abstract
High-level synthesis (HLS) is a widely used tool in designing Field Programmable Gate Array (FPGA). HLS enables FPGA design with software programming languages by compiling the source code into an FPGA circuit. The source code includes a program (called “kernel”) and several pragmas that instruct hardware synthesis, such as parallelization, pipeline, etc. While it is relatively easy for software developers to design the program, it heavily relies on hardware knowledge to design the pragmas, posing a big challenge for software developers. Recently, different machine learning algorithms, such as GNNs, have been proposed to automate the pragma design via performance prediction. However, when applying the trained model on new kernels, the significant domain shift often leads to unsatisfactory performance. We propose a more domain-generalizable model structure: a two-level hierarchical Mixture of Experts (MoE), that can be flexibly adapted to any GNN model. Different expert networks can learn to deal with different regions in the representation space, and they can utilize similar patterns between the old kernels and new kernels. In the low-level MoE, we apply MoE on three natural granularities of a program: node, basic block, and graph. The high-level MoE learns to aggregate the three granularities for the final decision. To train the hierarchical MoE stably, we further propose a two-stage training method to avoid expert polarization. Extensive experiments verify the effectiveness of the proposed hierarchical MoE. We publicized our codes at https://github.com/weikai-li/HierarchicalMoE.
@@ Introduction
With the heated demand for domain-specific accelerators, field-programmable gate arrays (FPGA) are widely used. It is, however, labor-intensive to write register-transfer-level hardware description languages, such as VHDL and Verilog. High-level synthesis (HLS) provides a much easier solution by compiling a source code written in a software programming language, such as C and MatLab, into an FPGA circuit (Cong et al. 2011, 2022; Schafer and Wang 2019).
Figure 1: Motivation of utilizing MoE. The two kernels share a similar nested loop, while the other parts are different. Similar parts can utilize similar experts via similar gating, while different parts can utilize different experts. This is just an illustration. Actually, we do not apply MoE on different granularities in the same model in this way. We aggregate them like Figure 2.
@@
While machine learning models have better learning ability than the heuristic methods, they do not generalize well to unseen kernels. In real-world applications, we often encounter new circuit design requirements. The model trained on existing kernels usually fails to perform well on new kernels. This can be viewed as a domain generalization (Bai et al. 2022; Kwon and Carloni 2020) problem, where each kernel is a domain. Note it is very time-consuming to run HLS to acquire the labels on new kernels, with each run taking minutes to hours (Sohrabizadeh et al. 2023), thus data-efficient fine-tuning is required. Domain generalization methods for GNN usually employ adversarial training to align the representation space between different domains (Dai et al. 2019; Zhang et al. 2019; Wu et al. 2020; Shen et al. 2023), or do data augmentation to learn invariant representation under risk extrapolation (Wu et al. 2022b; Liu et al. 2023a). However, these approaches are not applicable to the HLS prediction task, as the differences between kernels are informative for the prediction, so forcing them to the same distribution does not work.
@@
Nonetheless, a unique opportunity of the HLS prediction task is that kernels usually share some similar substructures, as shown in Fig. 1. It could be beneficial if we can leverage such similarities. We thus propose a two-level hierarchical mixture of experts (MoE) to address this issue. MoE (Jacobs et al. 1991; Shazeer et al. 2017) uses several expert networks instead of a single fixed network, and it computes the weighted sum of their outputs. Different experts can specialize in different regions in the hidden representation space, so that similar substructures in different domains can be processed by the same experts. A gating network computes the expert assignment, which learns which expert should deal with which substructure. GraphMETRO (Wu et al. 2024) applies MoE to improve GNNs’ generalizability by predefining a set of domain shifts and training each expert to deal with one type of domain shift by data augmentation. We take an orthogonal direction, eliminate the need for hardware knowledge to define the domain shifts, and focus on the model structure rather than data augmentation.
@@
Another unique opportunity for HLS prediction is that the input graph naturally has three levels of granularity: nodes, blocks, and the whole graph. MoE can be constructed on either of these three granularities. A natural question is which granularity to use. It is a hard task as different kernels benefit from different granularities. Simply applying MoE on all three granularities in one model does not work. Therefore, we propose a higher-level MoE to aggregate the three low-level MoEs according to the need. Nonetheless, optimization with the hierarchical MoE is challenging. Since the three low-level MoEs are different, training easily leads to expert polarization. We propose a two-stage training to stabilize the training. To the best of our knowledge, while some papers (Shazeer et al. 2017) have studied hierarchical MoE to improve memory efficiency, this is the first study that demonstrates the performance gain of hierarchical MoE. We conduct extensive experiments on the largest HLS benchmark dataset, HLSyn (Bai et al. 2023), and the experiment results reveal the effectiveness of hierarchical MoE. In summary, we make the following contributions:
@@
  • We propose hierarchical MoE to address these challenges and a two-stage training approach to stabilize its training.
@@
There are two directions for HLS pragma design automation. The first direction is heuristic-driven. AutoDSE (Sohrabizadeh et al. 2021) uses bottleneck-guided searching based on the HLS feedback, and it is well-generalizable. Non-linear programming performs well on affine kernels based on a lower-bound objective function (Pouget, Pouchet, and Cong 2024a, b). The second direction is data-driven, which trains a surrogate machine learning model to predict the FPGA’s latency and resource utilization. GNNs (Sohrabizadeh et al. 2022, 2023; Ustun et al. 2020; Wu et al. 2022a; Wu, Xie, and Hao 2023, 2021; Qin et al. 2024) are widely used. While machine learning models could perform better than heuristic methods with sufficient training data, they suffer from poor generalizability. Several previous works have tried to solve this issue. (Ferretti et al. 2020) transforms the best pragma design from the most similar source kernel to the target kernel, but its transformation is too simple and highly limited. (Kwon and Carloni 2020) trains a separate domain adaptor for each domain, while (Bai et al. 2022) uses the meta-learning method, MAML (Finn, Abbeel, and Levine 2017), to find a generalizable parameter initialization. However, since we have many kernels, it is difficult to train MAML to get stable results. We only have hundreds of data for each kernel, so it is difficult to train an adaptor. Instead, we explore an orthogonal direction: to improve the model structure’s generalizability.
@@
Domain generalization aims to reduce the performance gap between the source and target domains. Most related works (Dai et al. 2019; Zhang et al. 2019; Wu et al. 2020; Shen et al. 2023) employ adversarial training to align the representation space between the source and target domains. Some papers (Wu et al. 2022b; Liu et al. 2023a) design specific data augmentation methods and learn invariant representations under risk extrapolation. However, we cannot directly align the representation space of different kernels since their difference is large and is useful for the prediction. GraphMETRO (Wu et al. 2024) predefines several types of domain shifts and uses MoE to deal with them, where each expert deals with one type. We take an orthogonal direction and eliminate the need for hardware knowledge to define the types of domain shifts, and our model based on the three granularities is specifically designed for this task.
@@
MoE uses several expert networks and calculates the weighted sum of their outputs (Jacobs et al. 1991; Shazeer et al. 2017). It is useful in domain transfer learning in computer vision (Li et al. 2023; Zhong et al. 2023). It has also been utilized in GNN to improve the performance (Wang et al. 2023; Han et al. 2024), diversify node representations in fairness-augmented graphs (Liu et al. 2023b), and deal with class imbalance (Hu et al. 2021). Previous work (Shazeer et al. 2017) has employed hierarchical expert routing to improve memory efficiency, enabling a larger number of experts. However, previous papers have not shown the performance gain of hierarchical MoEs over regular MoEs when using the same number of total experts.
@@
HLS prediction is formalized as a graph regression task. Following previous works (Sohrabizadeh et al. 2022, 2023), we utilize ProGraML (Cummins et al. 2021) graph to represent a source code, which is a control data flow graph. Nodes represent instructions, variables, and constant values, and edges represent the control flow and data flow. A GNN model is trained to predict the FPGA’s latency and the utilization of four resources: LUT, FF, DSP, and BRAM. In the domain generalization setting, we train the model on $N$ source kernels $D^{(train)}=\{D_{1},D_{2},...,D_{N}\}$, where $D_{i}=\{(X_{i},T_{i1},Y_{i1}),(X_{i},T_{i2},Y_{i2}),...,(X_{i},T_{ii_{n}},Y_{ii_{n}})\}$ is the $i$-th kernel containing $i_{n}$ samples, and each sample consists of a program $X$, a pragma design $T$, and a label $Y$. For a new kernel $D_{test}$, we consider the constraint of data scarcity where we only use $k$ samples from the dataset $D^{(k)}\subset D_{test},|D^{(k)}|=k$, to fine-tune the model. After fine-tuning, there are two ways of evaluation: offline evaluation and online evaluation. In the offline evaluation, we evaluate the MSE on the left-out test samples: $D_{test}\setminus D^{(k)}$. In the online evaluation, we do design space exploration (DSE) to search for good pragma designs. Based on the fine-tuned model’s prediction, we select the top $M$ designs to run HLS, and evaluate the speedup ($M$ is set to 10 in our experiments).
@@
The proposed hierarchical MoE can be adapted to any GNN model. We use one of the SOTA GNN models for HLS prediction, HARP (Sohrabizadeh et al. 2023), as the base model. Since the original ProGraML graph is not good at modeling long-range dependency, HARP creates a “pseudo node” for each basic block and connects it with every node within that basic block. A basic block is a sequence of instructions with a single entry point and a single exit point where the terminator instruction can be a branch, return, etc. We illustrate HARP’s model structure in the appendix. We use $V$ to denote the set of all nodes and $V_{B}$ to denote the set of pseudo nodes. HARP consists of five components: (1) GNN encoder: it consists of 6 GNN layers and learns node representation; (2) Pragma MLP: it uses an MLP for each pragma type to transform the representation of pseudo nodes that are modified by that pragma type; (3) Another GNN layer after pragma MLP: it further updates the node representations; (4) Graph pooling: it performs graph pooling on pseudo nodes $V_{B}$ to form one graph representation and on normal nodes $V\setminus V_{B}$ to form another, then concatenates the two; (5) Output MLP: it makes the final prediction.
@@ Methods
A unique opportunity for the HLS prediction task is that the input graph has three granularities: normal nodes that represent data and instructions, pseudo nodes that represent basic blocks, and the whole graph that represents a source code file. If a data point from the target kernel shares a similar statement (node), basic block, or the whole code (graph) with a data point from the training kernels, MoE could be helpful. Thus, we consider employing MoE on the three granularities: MoE on all nodes including normal nodes and pseudo node (node MoE), MoE only on pseudo nodes (block MoE), and MoE on the graph (graph MoE). Based on our pre-exploration, the best practice is to apply MoE in the components after the pragma MLP, since the experts need to share the same encoder and pragma MLPs. In our pre-exploration, we found that the best structure of the gating network is a linear layer. The model structures are shown in Figure 2.
@@
$\boldsymbol{h_{i}^{\prime}}=\sum_{j=1}^{n}softmax(\boldsymbol{W_{G}^{node}}\boldsymbol{h_{i}})_{j}\cdot GNN_{j}(\boldsymbol{h_{i}},\{h_{j},j\in N(v_{i})\}),$ (1)
where $\boldsymbol{W_{G}^{node}}\in\mathbb{R}^{n\times d}$ is the parameter of the gating network, $N(v_{i})$ is the neighbors of $v_{i}$, and $GNN_{j}(\cdot)$ is the $j$-th expert. The $softmax$ function is a commonly-used normalization method that ensures the weights of all experts sum up to 1. The expert weights can be seen as “attention”. We calculate the expert weights (attention scores) using the gating network based on the node embeddings $\boldsymbol{h_{i}}$. By employing the MoE layer, different nodes can utilize different parameters for message passing and message aggregation.
Block MoE. We have tried applying MoE on the pragma MLP, but the performance has decreased. It might be because each pragma type has a separate pragma MLP, so the training data for each pragma MLP is already limited. If we apply MoE, it will dilute the training data even more. Instead, we add an additional MoE layer on the pseudo nodes’ representations after the pragma MLP, so that we can increase the expressiveness of the pseudo nodes’ representations before graph pooling. The representation space of the pseudo nodes after the block MoE could be different from normal nodes, so it is no longer suitable to have another GNN layer after the pragma MLP to do message-passing between pseudo nodes and normal nodes. We have also tried applying an additional layer for normal nodes after the pragma MLP, but the performance is unsatisfactory. It might be because the normal nodes lack a global view, so it is not useful to further transform its representation. After having the block MoE, performing graph pooling only on the pseudo nodes performs better than graph pooling on both pseudo nodes and normal nodes. Therefore, we do not utilize normal nodes after the GNN encoder in the block MoE model. Ablation study of various designs is in the appendix.
@@
$\boldsymbol{h_{i}^{\prime}}=\sum_{j=1}^{n}softmax(\boldsymbol{W_{G}^{block}}\boldsymbol{h_{i}})_{j}\cdot\boldsymbol{W_{j}}\boldsymbol{h_{i}}\ \ (v_{i}\in V_{B}),$ (2)
where $\boldsymbol{W_{G}^{block}}\in\mathbb{R}^{n\times d}$ is the parameter of the gating network, and $\boldsymbol{W_{j}}\in\mathbb{R}^{d\times d}$ is the $j$-th expert network.
@@
$\hat{Y}^{(t)}=\sum_{j=1}^{n}softmax(\boldsymbol{W_{G}^{graph}}\boldsymbol{h_{G}})_{j}\cdot MLP^{(t)}_{j}(\boldsymbol{h_{G}}),$ (3)
where $W_{G}^{graph}\in\mathbb{R}^{n\times d}$ is the parameter of the gating network, $MLP^{(t)}_{j}(\cdot)$ is the $j$-th expert for predicting the objective $t$ ($t$ could be latency or a certain resource’s utilization). The expert assignment is the same for all prediction objectives. The MLP contains four linear layers and the ELU activation function in between. Different design points can utilize different experts to make the final prediction.
Regularization term. As discovered in many literature (Shazeer et al. 2017; Wang et al. 2023; Li et al. 2023), expert polarization is a common problem of MoE. Due to the random initialization of expert networks, different experts initially perform differently. The gating network learns to assign higher weights to better experts. It results in more training of the initial better experts, leading to their even better performance, which forms a cycle. As training continues, the MoE model might collapse and only use the best expert. To avoid this issue, we apply a regularization term commonly used in MoE (Shazeer et al. 2017):
@@
where $CV(\cdot)$ is the coefficient of variation. We calculate the importance score $\boldsymbol{I(W_{G})}\in\mathbb{R}^{n}$ for gating network $W_{G}$ as the total weights assigned to each expert. For node/block MoE, $M$ is the total number of nodes/pseudo nodes in all graphs; for graph MoE, $M$ is the number of graphs. This regularization encourages balanced expert assignments.
@@
We propose two designs of the high-level gating network, as illustrated in Figure 2. The first design is to perform graph pooling on the input node features to form a graph representation as the input to the high-level gating network. We use the self-attention graph pooling. Denoting the input feature of node $v_{i}$ as $\boldsymbol{x_{i}}$, the graph pooling is formalized as:
@@
where $X_{G}$ is the aggregated input feature, and $V$ is the set of all nodes. The second design is to concatenate the graph representation in the three low-level MoE models. The second design performs better in our experiments, since the hidden representations are more expressive than the input features. Nonetheless, when we use a sparse MoE where only one or two experts are selected, the first design will be more memory efficient. It can determine the expert assignment before the computation of three expert models, thus reducing unnecessary computation. However, the best-performing method is to utilize all the experts, and in this case, the two designs are similar in efficiency. We utilize all the experts in our main experiments. We also apply the regularization term for the high-level gating network.
Two-Stage Training Strategy
Optimizing the hierarchical MoE is challenging. Different from previous MoE studies where expert networks have the same structures, our three low-level MoEs are very different and thus have different convergence speeds. The graph MoE model converges the fastest, since its MoE operates on the graph representation and has the least computation. Thus, the high-level gating network suffers more severely from expert polarization. It learns to assign nearly all the weights to the graph MoE model. If we simply increase the weight of the regularization term, the high-level gating network will learn to assign about $\frac{1}{3}$ weight to each expert, but the graph MoE model still converges the fastest. As a result, the graph MoE model will learn to output three times the prediction, while node and block MoE models will learn to output zero.
To address this unique challenge, we design a two-stage training strategy to encourage every expert model to perform well. In the first stage containing $T$ epochs (warmup), we train the three expert models individually. In the second stage, we take turns training the whole model end-to-end and the three expert models individually. If we denote the label as $Y$, the prediction made by the $i$-th expert model as $\hat{Y}_{i}$, and the MSE loss function as $L(Y,\hat{Y_{i}})$, then we define the loss function at epoch $t$ as:
@@
Here, $L_{R}=\frac{1}{3}(L_{R1}+L_{R2}+L_{R3})$, where $L_{Ri}$ is the regularization term of the i-th low-level MoE. $L_{Rh}$ is the regularization term of the high-level MoE. $g_{i}$ is the weight assigned by the high-level gating network.
This two-stage training strategy does not increase the training time and is easy to implement. We only need to disable the high-level gating network and change the loss function in certain epochs. This strategy is only used during training. When we fine-tune a trained model on target kernels, we directly use the end-to-end joint training, since all the experts can already perform well and there is no risk of polarization. Besides, we initialize the high-level gating network to assign the same weights to the expert models, which can further prevent expert polarization. For low-level MoEs, we use the normal random initialization.